OpenStudy (anonymous):

Simplify: Negative two and one-third minus negative ten and one-sixth Negative twelve and one-half Twelve and one-half Seven and five-sixths Negative seven and five-sixths

OpenStudy (anonymous):

(62+(1/3)) - (10+(1/6)) -------------------- (1/3) = (2/6) (62+(2/6)) - (10+(1/6)) = (62-10) - ((2/6) - (1/6)) = 52 and (1/6)